Circular No. 9069

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2009 (93) 1 AND S/2009 (93) 2 F. Marchis and B. Macomber, Carl Sagan Center at the SETI Insti-

tute and University of California at Berkeley; J. Berthier and F. Vachier, Institut de Mécanique Céleste et de Calcul des Ephémérides, Observatoire de Paris; and J. P. Emery, University of Tennessee, Knoxville, report that adaptive-optics images were recorded of (93) Minerva, a large C-type mainbelt asteroid, with the 10-m Keck II telescope [+ NIRC2 camera (angular resolution up to 0′′.042, corresponding to a spatial resolution of 65 km) + Fe II filter (central wavelength at 1.64 μm)] between Aug. 16.57 and 16.64 UT, with the target at r = 2.117 AU and phase angle 20^{o}.0. The direct images reveal that the large minor planet has an almost-spherical shape with an average diameter of 145 km. Additionally, these observations show the presence of a companion ∼ 4 km in diameter at 0′′.41 (projected distance of 630 km) in p.a. 275^{o} detected in every image recorded over the 1.8-hr baseline. Careful analysis of three images indicates the presence of a second satellite (∼ 3 km) and located closer to the primary at an apparent distance of 380 km (0′′.25) in p.a. 209^{o}. With (87) Sylvia (cf. IAUC 8582),

(45) Eugenia (cf. IAUC 8817), and (216) Kleopatra (cf. IAUC 8980), this is the fourth multiple system with two small-kilometer-sized satellites to be discovered and imaged in the main asteroid belt.

COMET C/2009 Q4 (BOATTINI) A. Boattini reports his discovery of a strongly condensed comet with

coma diameter ∼ 8′′ (slightly elongated in p.a. 250^{o}) in CCD images taken with the Catalina 0.68-m Schmidt telescope (discovery observation tabulated below); Catalina images taken by R. E. Hill on Aug. 28.4 UT show a small nuclear condensation with a 2′′ coma and a narrow tail ∼ 10′′–15′′ in p.a. 290^{o}. Following posting on the ‘NEOCP’ webpage, R. Ligustri (Udine, Italy) writes that his CCD images on Aug. 29.4 (remotely using a 0.25-m reflector near Mayhill, NM) show a slightly diffuse coma with diameter ∼ 15′′. J. V. Scotti notes that images taken on Aug. 30.5 with the Spacewatch 1.8-m reflector show a 9′′ coma and a 0′.68 tail in p.a. 257^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Aug. 26.47241 4^{h}27^{m}20.^{s}05 +8^{o}18′42′′.1 18.9 Boattini

The available astrometry, preliminary parabolic orbital elements [T = 2009 Nov. 1.603 TT, q = 1.50024 AU, i = 11^{o}.627, ω = 301^{o}.086, Ω = 135^{o}.512 (equinox 2000.0)], and an ephemeris appear on MPEC 2009-Q84.

2009 August 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green